OpenStudy (anonymous):

Find the zeros of the polynomial function and state the multiplicity of each. f(x) = 5(x + 7)^2(x - 7)^3

OpenStudy (mathstudent55):

Set each factor containing a variable equal to zero. Then solve for x. The multiplicity of each zero is the exponent of the factor that gave you that zero.

OpenStudy (agent0smith):

\[\Large 5(x + 7)^2(x - 7)^3 = 0\]means either x+7=0 or x-7=0. The power 2 on (x+7) means it is a doubled zero. The power 3 on (x-7) means it's a tripled zero.

OpenStudy (anonymous):

Long story short. You have two factors: (x + 7) and (x - 7) Those give you the zeros like so: x + 7=0 x = -7 and x - 7 = 0 x = 7 The zeros is the hardest part. Multiplicity can be found by the exponent of the factors. (x + 7)² This is to the second power. This is the same as: (x + 7)(x + 7) So the multiplicity is 2. (x - 7)³ This is to the third power. This is the same as: (x - 7)(x - 7)(x - 7). So the multiplicity is 3. Make sense?

OpenStudy (anonymous):

the multiplicites are your exponents basically like if i said (x-1)^3 your root is 1 and your multiplicites are 3 or is that 2 since you already counted the 1 once as a root? anyhoo that is cuzz (x-1)^3=(x-1) (x-1) (x-1) so the root 1 accurs 3 times.

OpenStudy (anonymous):

5(x + 7)^2(x - 7)^3 Usually the outside numbers dont matter so you can have (x + 7)^2(x - 7)^3 x=-7 and x=7 with like 5 multiplicites. I forget if you include the root or not. If not then it is 3

OpenStudy (anonymous):

-7, multiplicity 2; 7, multiplicity 3 4, multiplicity 1; -7, multiplicity 3; 7, multiplicity 3 -7, multiplicity 3; 7, multiplicity 2 4, multiplicity 1; 7, multiplicity 1; -7, multiplicity 1

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

top one ?

OpenStudy (cggurumanjunath):

\[5(x+7)^{2}(x-7)^{3}\]

OpenStudy (anonymous):

?

OpenStudy (ybarrap):

http://tinyurl.com/lornbeach1

OpenStudy (ybarrap):

Solve for x over the real numbers: \(5 (x-7)^3 (x+7)^2 = 0 \) Divide both sides by 5: \((x-7)^3 (x+7)^2 = 0\) Split into two equations: \((x-7)^3 = 0 or (x+7)^2 = 0\) Take cube roots of both sides: \(x-7 = 0\) or \((x+7)^2 = 0\) Add 7 to both sides: \(x = 7\) or\) \((x+7)^2 = 0\) Take the square root of both sides: \(x = 7 \)or \(x+7 = 0\) Subtract 7 from both sides: so \( x = 7\) or \(x = -7\)

OpenStudy (anonymous):

for anyone needing the answer to this it's: -7 with a multiplicity of 2 and 7 with a multiplicity of 3 :)